# Quasi-continuous sub-K strontium source without a high-finesse cavity stabilized laser

### v1

@@ I Introduction
To overcome these limitations, we introduce a dispersion-engineered frequency comb and demonstrate its use for a quasi-continuous ultracold strontium source. The frequency comb is generated using mode-locked Er:fiber oscillators, which are intrinsically insensitive to pump-induced fluctuations by dispersion engineering. These measures preserve the coherence and spectral purity of individual comb teeth, enabling direct laser stabilization with sub-kHz precision by locking its repetition-rate to an RF reference. The reduced susceptibility to thermal and vibrational noise also enables easier integration into compact, transportable quantum platforms. Through a non-linear conversion processes, a broad spectral range extending from 420 nm to beyond 2000 nm can be obtained, creating an architecture adaptable for narrow-linewidth laser cooling of strontium at 429 THz but also for a wide range of other narrow-linewidth alkaline-earth(-like) [26, 27] and lanthanide atoms [28, 29]. We use our fully fiber-based frequency comb to generate the ultracold strontium source by directly locking the narrow-line cooling laser to it without any need of a high-finesse free-space reference cavity. Thanks to its broad coverage and flexible offset-locking capabilities, our system supports laser cooling of all three stable bosonic strontium isotopes using a single unified laser architecture. Combined with a robust laser cooling setup, this enables the generation of a continuous source of ultra-cold strontium atoms at K temperatures. To ensure long-term frequency stability of the comb, its repetition rate must be stabilized to an external reference. We employ two approaches, which we compare against a cavity-based laser system: a highly stable external RF reference, specifically a 10 MHz signal provided by the Dutch Metrology Institute VSL [30], and a self-contained stabilization scheme locking the comb to a tunable oven-controlled crystal oscillator, which is referenced to the MOT position [2]. The scheme enables robust operation even in the absence of external frequency standards. Finally, we introduce a technique for quasi-continuous out-coupling of sub-K atoms from the MOT, highlighting the potential of this platform for zero-dead-time optical lattice clocks [31, 32], portable clocks [33, 34], and quantum devices that require quasi-continuous atomic sources [15, 35, 9, 16, 17].
@@ II.2 SHORT-TERM STABILITY
As shown in Figure 2, with increasing pump power, the fixed point frequency shifts considerably, reaching a value close to the target atomic transition frequency of 429 THz at approximately 161 mW. The measured linewidth data shows a clear correlation with the fixed point behavior. At pump powers far from 161 mW, the linewidth is substantially broadened, indicating that noise contributions from the pump strongly affect the spectral purity. However, near 161 mW, the linewidth reaches a minimum below 1 kHz, suggesting that the effective noise at the transition frequency is minimized due to the proximity of the fixed point. By selecting a pump power where the fixed point aligns with the atomic transition frequency, high spectral purity and long-term stability can be obtained without the complexity of optical cavity locking.
@@ II.3 LONG-TERM STABILITY
However, locking and is essential to stabilize the frequency comb. Any significant drift away from the optimal pump power alters the intra-cavity conditions, shifts the pump-induced fixed point, and increases noise at the target atomic transition. To address this, we ensure that the stabilization scheme maintains the pump power within a narrow window around the optimal point (e.g., 161 mW), where the fixed point aligns with 429 THz and sub-kHz linewidths are achieved.
@@ III.1 CONTINUOUS ULTRACOLD STRONTIUM SOURCE
The design of our continuous cold atom apparatus is derived from previous work in our group [12, 9]. We have simplified the previous design, reducing the total volume of the vacuum system to just m3. The continuous ultracold atom source is created through the simultaneous execution of a sequence of cooling stages that are separated in space, as shown in Fig. 3(a). The sequence starts with a hot atomic beam that effuses out of an oven heated to approximately . The atomic beam is cooled along the two transverse axes using two retroreflected beams addressing the blue cooling transition (1SP1, see Fig. 3(b)). We slow the atomic beam with a permanent-magnet Zeeman slower using mW of light 242 MHz red-detuned from the blue transition. The slowed atoms are captured by a 2D magneto-optical trap (MOT) operating on the broad blue cooling transition and cooled to temperatures on the order of a few mK. The quadrupole magnetic field for the 2D MOT is produced by stacks of permanent magnets and has a gradient of 20 G/cm. The 2D MOT laser beams are 32 MHz red-detuned from the blue cooling transition and have a beam waist of 1.5 cm and a peak intensity of 0.4 . Since there is no confinement in the direction of gravity in the 2D MOT, atoms can fall freely out of the MOT and into a lower chamber. The atoms’ upward movement is limited by two 461-nm “push” beams angled symmetrically about the -axis. As the atoms drop to the bottom chamber, the atomic beam is collimated directly below the blue 2D MOT beams, using two red molasses beams addressing the narrow red cooling transition (1SP1, 7.5 kHz wide). The average time that atoms spend in the blue MOT is less than 5 ms, enabling high atomic flux into the lower chamber without the need for repump lasers. This is possible despite the partially open nature of the 1SP1 transition, which allows decay into metastable states [22, 46].
@@
As atoms arrive in the bottom chamber containing a broadband five-beam red MOT (BB MOT), comprised of two counterpropagating beam pairs along the horizontal axes and one upwards propagating MOT beam. The latter beam is modulated at a frequency of 70 kHz over a range of MHz red-detuned from the MOT transition, allowing it to also act as a white light slower [47]. The horizontal MOT beams are retro-reflected, with a waist of 15 mm and peak intensity of 1.1 mW/cm2. They are modulated at a frequency of 30 kHz to cover a range of MHz red-detuned from the MOT transition. The quadrupole magnetic field is produced by a quadrupole coil pair, with a gradient of approximately G/cm. Notably, only five laser beams are used to produce the red MOT. There is no need for a vertically downwards propagating MOT beam, as gravity provides a sufficient force to balance out the force of the upwards propagating MOT beam. With these parameters, we achieve a steady state BB red MOT of 88Sr with atoms. For further information on the apparatus design, see Ref. [48].
@@ III.2 DIFFERENT REFERENCES FOR LONG-TERM STABILITY
For the first method, we use a 10 MHz RF source from the Dutch Metrology Institute (VSL) as a reference to the frequency comb. This RF signal is produced by a local White Rabbit switch locked to a White Rabbit grandmaster (GM) switch located at VSL Delft, which itself is referenced to UTC(VSL). The White Rabbit switches are connected through about 100 km of optical fiber of the SURF research and education network. The White Rabbit link forms part of a larger network, the SURF Time&Frequency network, which is used to disseminate UTC(VSL) to multiple locations in the Netherlands via the White Rabbit protocol. Typical frequency stability of the provided 10 MHz is between and fractionally after 1 s of averaging, dropping down to after 100 s of averaging, and dominated by white phase noise between 1 and 100 s of averaging [49, 50, 51, 30].
@@
We examine the stability of the BB MOT position using the conventional laser system and using the comb laser system with each of the two comb locking methods over a period of 900 s each, as shown in Fig. 4. While the conventional laser system provides the best short-term stability, with a standard deviation of the position of 27.1 m over 900 s, there are larger variations in the mean MOT position over longer time scales, but no clear drift. With the comb locked to the VSL RF source, the standard deviation of the MOT position over 900 s is increased to 60.73 m. A steady drift is observed over the whole measurement period, which could come from a drift of the VSL RF frequency or of the comb repetition rate locking electronics. Other reasons could be magnetic field or MOT laser intensity drift. However, those drifts would also influence the MOT position when using the cavity-locked laser system, which we don’t observe. Stabilizing the comb on the red MOT position produces a standard MOT position deviation of 83.6 m over 900 s. Long-term drift is absent as it is removed by the MOT-position lock. If the MOT fluorescence imaging system is stable, this can contribute to more reliable loading of MOT atoms into dipole traps compared to the previous two methods, rendering subsequent experimental stages more reliable. We also consider the standard deviation of the MOT position over a 2 s interval: approximately 10 m for the conventional laser system, 35 m for the VSL, and 85 m for the MOT position stabilization scheme. For the comb-locked methods, the stability over 2 s is primarily limited by the stability of the RF source used to synchronize the frequency comb within the same time frame. In particular, the performance of the MOT position stabilization scheme could be significantly improved by employing a quartz oscillator with enhanced short-term stability.
@@
We also characterize a single-frequency (SF) MOT obtained using the three different laser locking methods. We create the SF MOT in a pulsed manner by ramping the range of the BB modulation of the MOT beams to zero, sweeping the center frequency to 700 kHz red-detuned from resonance, and simultaneously reducing the beam intensity to about 250 W/cm2. We then further ramp the intensity down to 20 W/cm2. During this process, the magnetic field gradient remains at 0.39 G/cm. This procedure allows atoms to be cooled to sub-K temperatures. Table 2 lists the temperatures achieved in the BB and SF MOTs with the conventional laser system and the comb based system referenced using either method, measured using time of flight [52]. We achieve temperatures below 10 K in the BB MOT using all locking methods. While the cavity-locked system and the comb-based system with MOT-position stabilization produce MOTs with very similar temperatures, we believe the comb-based system referenced to VSL produces a slightly higher temperature due to reduced polarization purity of the MOT beams when this data set was taken. In all cases, we are able to achieve sub-K temperatures in the SF MOT with around atoms, and we see comparable performance between the cavity-stabilized system and the comb-stabilized system. SF MOT temperatures depend on MOT density, MOT laser intensity and the short-term ( ms) red MOT laser linewidth. The lowest temperature can only be achieved if the laser linewidth is narrower than the red MOT transition’s natural linewidth of 7.5 kHz. The lowest temperatures that have been reached are around 0.4 K for MOT laser intensities around [18]. In all three cases we reach temperatures that are less then twice as large, on par with expectation for our MOT laser intensity of several , indicating a sufficiently narrow linewidth of the MOT laser.
@@ Acknowledgements.
We thank Rob Smets for initiating the first White Rabbit implementations and laying the foundation for the White Rabbit service at SURF. We thank OPNT B.V. for setting up and calibrating the White Rabbit network that was used during our research. We thank N.J. van Druten, S. Wolzak and S. Sarkar for establishing the White Rabbit connection at UvA. We thank N.J. van Druten for thoughtful discussions and careful reading of the manuscript. This work has received funding from the European Union’s (EU) Horizon 2020 research and innovation program under Grant Agreement No. 820404 (iqClock project) and No. 860579 (MoSaiQ project). It also received funding from the European Partnership on Metrology, co-financed from the European Union’s Horizon Europe Research and Innovation Programme and by the Participating States, under project 22IEM01 TOCK. It further received funding from the Dutch National Growth Fund (NGF), as part of the Quantum Delta NL programme.Appendix A DUAL MOT LIFE TIME, TEMPERATURE AND LONG TERM STABILIZATION SCHEME

### v2

@@ I Introduction
To overcome these limitations, we introduce a dispersion-engineered frequency comb and demonstrate its use for a quasi-continuous ultracold strontium source. The frequency comb is generated using mode-locked Er:fiber oscillators, which are intrinsically insensitive to pump-induced fluctuations by dispersion engineering. These measures preserve the coherence and spectral purity of individual comb teeth, enabling direct laser stabilization with sub-kHz precision by locking its repetition-rate to an RF reference. The reduced susceptibility to thermal and vibrational noise also enables easier integration into compact, transportable quantum platforms. Through a non-linear conversion processes, a broad spectral range extending from 420 nm to beyond 2000 nm can be obtained, creating an architecture adaptable for narrow-linewidth laser cooling of strontium at 435 THz but also for a wide range of other narrow-linewidth alkaline-earth(-like) [26, 27] and lanthanide atoms [28, 29]. We use our fully fiber-based frequency comb to generate the ultracold strontium source by directly locking the narrow-line cooling laser to it without any need of a high-finesse free-space reference cavity. Thanks to its broad coverage and flexible offset-locking capabilities, our system supports laser cooling of all three stable bosonic strontium isotopes using a single unified laser architecture. Combined with a robust laser cooling setup, this enables the generation of a continuous source of ultra-cold strontium atoms at K temperatures. To ensure long-term frequency stability of the comb, its repetition rate must be stabilized to an external reference. We employ two approaches, which we compare against a cavity-based laser system: a highly stable external RF reference, specifically a 10 MHz signal provided by the Dutch Metrology Institute VSL [30], and a self-contained stabilization scheme locking the comb to a tunable oven-controlled crystal oscillator, which is referenced to the MOT position [2]. The scheme enables robust operation even in the absence of external frequency standards. Finally, we introduce a technique for quasi-continuous out-coupling of sub-K atoms from the MOT, highlighting the potential of this platform for zero-dead-time optical lattice clocks [31, 32], portable clocks [33, 34], and quantum devices that require quasi-continuous atomic sources [15, 35, 9, 16, 17].
@@ II.2 SHORT-TERM STABILITY
As shown in Figure 2, with increasing pump power, the fixed point frequency shifts considerably, reaching a value close to the target atomic transition frequency of 435 THz at approximately 161 mW. The measured linewidth data shows a clear correlation with the fixed point behavior. At pump powers far from 161 mW, the linewidth is substantially broadened, indicating that noise contributions from the pump strongly affect the spectral purity. However, near 161 mW, the linewidth reaches a minimum below 1 kHz, suggesting that the effective noise at the transition frequency is minimized due to the proximity of the fixed point. By selecting a pump power where the fixed point aligns with the atomic transition frequency, high spectral purity and long-term stability can be obtained without the complexity of optical cavity locking.
@@ II.3 LONG-TERM STABILITY
However, locking and is essential to stabilize the frequency comb. Any significant drift away from the optimal pump power alters the intra-cavity conditions, shifts the pump-induced fixed point, and increases noise at the target atomic transition. To address this, we ensure that the stabilization scheme maintains the pump power within a narrow window around the optimal point (e.g., 161 mW), where the fixed point aligns with 435 THz and sub-kHz linewidths are achieved.
@@ III.1 CONTINUOUS ULTRACOLD STRONTIUM SOURCE
The design of our continuous cold atom apparatus is derived from previous work in our group [12, 9]. We have simplified the previous design, reducing the total volume of the vacuum system to just m3. The continuous ultracold atom source is created through the simultaneous execution of a sequence of cooling stages that are separated in space, as shown in Fig. 3(a). The sequence starts with a hot atomic beam that effuses out of an oven heated to approximately . The atomic beam is cooled along the two transverse axes using two retroreflected beams addressing the blue cooling transition (1SP1, see Fig. 3(b)). We slow the atomic beam with a permanent-magnet Zeeman slower using mW of light 242 MHz red-detuned from the blue transition. The slowed atoms are captured by a 2D magneto-optical trap (MOT) operating on the broad blue cooling transition and cooled to temperatures on the order of a few mK. The quadrupole magnetic field for the 2D MOT is produced by stacks of permanent magnets and has a gradient of 20 G/cm. The 2D MOT laser beams are 32 MHz red-detuned from the blue cooling transition and have a beam waist of 1.5 cm and a peak intensity , where is the saturation intensity of the blue cooling transition. Since there is no confinement in the direction of gravity in the 2D MOT, atoms can fall freely out of the MOT and into a lower chamber. The atoms’ upward movement is limited by two 461-nm “push” beams angled symmetrically about the -axis. All laser beams addressing the blue cooling transition are produced by injection-locked lasers, for which the parent laser is stabilized using modulation transfer spectroscopy from a hot Sr vapor. As the atoms drop to the bottom chamber, the atomic beam is collimated directly below the blue 2D MOT beams, using two red molasses beams addressing the narrow red cooling transition (1SP1, 7.5 kHz wide). The average time that atoms spend in the blue MOT is less than 5 ms, enabling high atomic flux into the lower chamber without the need for repump lasers. This is possible despite the partially open nature of the 1SP1 transition, which allows decay into metastable states [22, 46].
@@
As atoms arrive in the bottom chamber containing a broadband five-beam red MOT (BB MOT), comprised of two counterpropagating beam pairs along the horizontal axes and one upwards propagating MOT beam. The latter beam is modulated at a frequency of 70 kHz over a range of MHz red-detuned from the MOT transition, allowing it to also act as a white light slower [47]. The horizontal MOT beams are retro-reflected, with a waist of 15 mm and peak intensity of 1.1 mW/cm2, or . They are modulated at a frequency of 30 kHz to cover a range of MHz red-detuned from the MOT transition. The quadrupole magnetic field is produced by a quadrupole coil pair, with a gradient of approximately G/cm. Notably, only five laser beams are used to produce the red MOT. There is no need for a vertically downwards propagating MOT beam, as gravity provides a sufficient force to balance out the force of the upwards propagating MOT beam. With these parameters, we achieve a steady state BB red MOT of 88Sr with atoms. For further information on the apparatus design, see Ref. [48].
@@ III.2 DIFFERENT REFERENCES FOR LONG-TERM STABILITY
For the first method, we use a 10 MHz RF source from the Dutch Metrology Institute (VSL) as a reference to the frequency comb. This RF signal is produced by a local White Rabbit switch locked to a White Rabbit grandmaster (GM) switch located at VSL Delft, which itself is referenced to UTC(VSL). The White Rabbit switches are connected through about 100 km of optical fiber of the SURF research and education network. The White Rabbit link forms part of a larger network, the SURF Time&Frequency network, which is used to disseminate UTC(VSL) to multiple locations in the Netherlands via the White Rabbit protocol. Typical frequency stability of the provided 10 MHz is between and fractionally after 1 s of averaging, dropping down to after 100 s of averaging, and dominated by white phase noise between 1 and 100 s of averaging [49, 50, 51, 30, 52].
@@
We examine the stability of the BB MOT position using the conventional laser system and using the comb laser system with each of the two comb locking methods over a period of 900 s each, as shown in Fig. 4. While the conventional laser system provides the best short-term stability, with a standard deviation of the position of 27.1 m over 900 s, there are larger variations in the mean MOT position over longer time scales, but no clear drift. With the comb locked to the VSL RF source, the standard deviation of the MOT position over 900 s is increased to 60.73 m. A steady drift is observed over the whole measurement period, which could come from a drift of the VSL RF frequency or of the comb repetition rate locking electronics. Other reasons could be magnetic field or MOT laser intensity drift. However, those drifts would also influence the MOT position when using the cavity-locked laser system, which we don’t observe. Stabilizing the comb on the red MOT position produces a standard MOT position deviation of 83.6 m over 900 s. Long-term drift is absent as it is removed by the MOT-position lock. If the MOT fluorescence imaging system is stable, this can contribute to more reliable loading of MOT atoms into dipole traps compared to the previous two methods, rendering subsequent experimental stages more reliable. We further analyze the short-term stability by dividing the data into 400 non-overlapping intervals of 2 s duration and computing the standard deviation of the MOT position within each interval. The reported values correspond to the mean of these interval-wise standard deviations, with the uncertainty given by their standard deviation. For the conventional laser system, we obtain , for the VSL , and for the MOT position stabilization scheme . For the comb-locked methods, the stability over 2 s is primarily limited by the stability of the RF source used to synchronize the frequency comb within the same time frame. In particular, the performance of the MOT position stabilization scheme could be significantly improved by employing a quartz oscillator with enhanced short-term stability.
@@
We also characterize a single-frequency (SF) MOT obtained using the three different laser locking methods. We create the SF MOT in a pulsed manner by ramping the range of the BB modulation of the MOT beams to zero, sweeping the center frequency to 700 kHz red-detuned from resonance, and simultaneously reducing the beam intensity to about 250 W/cm2. We then further ramp the intensity down to 20 W/cm2. During this process, the magnetic field gradient remains at 0.39 G/cm. This procedure allows atoms to be cooled to sub-K temperatures. Table 2 lists the temperatures achieved in the BB and SF MOTs with the conventional laser system and the comb based system referenced using either method, measured using time of flight [53]. We achieve temperatures below 10 K in the BB MOT using all locking methods. While the cavity-locked system and the comb-based system with MOT-position stabilization produce MOTs with very similar temperatures, we believe the comb-based system referenced to VSL produces a slightly higher temperature due to reduced polarization purity of the MOT beams when this data set was taken. In all cases, we are able to achieve sub-K temperatures in the SF MOT with around atoms, and we see comparable performance between the cavity-stabilized system and the comb-stabilized system. SF MOT temperatures depend on MOT density, MOT laser intensity and the short-term ( ms) red MOT laser linewidth. The lowest temperature can only be achieved if the laser linewidth is narrower than the red MOT transition’s natural linewidth of 7.5 kHz. The lowest temperatures that have been reached are around 0.4 K for MOT laser intensities around [18]. In all three cases we reach temperatures that are less than twice as large, on par with expectation for our MOT laser intensity of several , indicating a sufficiently narrow linewidth of the MOT laser.
@@ Acknowledgements.
We thank Rob Smets for initiating the first White Rabbit implementations and laying the foundation for the White Rabbit service at SURF. We thank OPNT B.V. for setting up and calibrating the White Rabbit network that was used during our research. We thank N.J. van Druten, S. Wolzak and S. Sarkar for establishing the White Rabbit connection at UvA. We thank N.J. van Druten for thoughtful discussions and careful reading of the manuscript. This work has received funding from the European Union’s (EU) Horizon 2020 research and innovation program under Grant Agreement No. 820404 (iqClock project) and No. 860579 (MoSaiQ project). It also received funding from the European Partnership on Metrology, co-financed from the European Union’s Horizon Europe Research and Innovation Programme and by the Participating States, under project 22IEM01 TOCK. It further received funding from the Dutch National Growth Fund (NGF), as part of the Quantum Delta NL programme. The raw data and the analysis tools used in this manuscript can be found in Reference [56].Appendix A DUAL MOT LIFE TIME, TEMPERATURE AND LONG TERM STABILIZATION SCHEME